Find the solution of the recurrence relation a_n=4a_(n-1)-4a_(n-2)+(n+1).2^n.
1
Expert's answer
2020-06-08T20:33:02-0400

Given an=4a(n-1)-4a(n-2)+(n+1).2n

Roots Characteristics equation

Let an=r2,a(n-1)=r and a(n-2)=1

r2=4r-4

r2 -4r+4=0

(r-2)2=0

r=2 with multiciplity 2

solution homogeneous linear recurrence relation

an=a1r1n+a2nr1n+------+akn(k-1)r1n with r1 a root with multiplicity k of the characteristic equation.

an(h)=a1 2n+ a2n2n

Particular solution

F(n)=(n+1)2n

Since 2 is a root of the characteristic equation with multiplicity 2

So particular solution takes the forms of

an(p)=n2(p1n+p0)2n

Thus the complete solution is the sum of homogeneous solution and particular integral

an=an(h)+an(p)=a12n+ a2n2n+n2(p1n+p0)2n

In order to find out the coefficient , P.I needs to satisfy the recurrence relation:

an=4(a-1)-4(a-2)+(n+1)2n

p1n3+p0n22n=4p1(n-1)32(n-1)+4p0(n-1)22n-1-4p1(n-2)32n-2-4p0(n-2)22n-2+(n+1)2n

4p1n3+4p0n2=8p1(n-1)3+8p0(n-1)2-4p1(n-2)3-4p1(n-2)3-4p0(n-2)2+4(n+1)

4p1n3+4p0n2=p1(4n3-24n+24)+p0(4n2-8)+4n+4

0=(-24p1+4)n+(24p1-8p0+4)

Equate the coefficients

0=(-24p1+4) and (24p1-8p0+4)=0

Thus p1=1/6 and p0=(24p1+4)/8=(24(1/6)+4)/8=1

Put the value of coefficients in complete solution

we have,

an=an(h)+an(p)=a12n+ a2n2n+n2(n/6+1)2n

an=(a1+a2n+n2+n3/6)2n